OpenStudy (anonymous):

Suppose r, s are elements of the real numbers. Prove the following: 1. r is an element of the rational numbers and (r+s) is not rational implies s is not rational 2. r is not rational, s is not equal to 0, and rs is rational implies s is not rational

ganeshie8 (ganeshie8):

prove the contrapositive for #1 : if s and r are rational, then r+s is rational

OpenStudy (anonymous):

I thought the contrapositive would be claiming: If s is rational \(\implies\) (r+s) is rational and r is not rational. I thought the implication flipped and all the statements were negated.

OpenStudy (freckles):

I thought it was If s is rational, then r is not rational or (r+s) is rational.

OpenStudy (freckles):

you know or instead of and

OpenStudy (anonymous):

I guess it would make sense to know which is which first, haha

ganeshie8 (ganeshie8):

oops! yeah \(\large \neg(A\land B) = \neg A \lor \neg B\)

OpenStudy (anonymous):

Okay, that makes more sense. But yeah, so if I list it out: A = r is rational B = r+s is not rational C = s is not rational. So I would want the contrapositive of (a^b) implies c, which is.... C v not(A^B) Yeah, I forgot how to properly negate an implication, whoops x_x

OpenStudy (freckles):

I think you can prove the first one by contradiction.

OpenStudy (anonymous):

i agree, contradiction seems natural to me in this case.

OpenStudy (freckles):

\[\text{ Assume } s=\frac{a}{b} , a,b \in \mathbb{Z} \text{ with } b \neq 0\]

OpenStudy (anonymous):

Yeah, I wasn't sure which method worked. I think I misunderstood contrapositive, but either way I thought it looked bad. So, which part of the statement would I contradict? Is there a specific part of the statement that must be chosen for the contradiction, or is it up to me really? Like, yeah, I think the natural one would be to say s = a/b, but is that the choice that must be made for contradiction? Just a conceptual question.

OpenStudy (anonymous):

If you want to prove \(A\Longrightarrow B\) by contradiction, you have to show that \(A\text{ and } \neg B\) leads to a contradiction. You don't get to choose what to negate. You have to negate the conclusion.

OpenStudy (anonymous):

Even though there are 3 statements, we're counting the first two as being A?

OpenStudy (anonymous):

Yep.

ganeshie8 (ganeshie8):

first two are premesis

OpenStudy (anonymous):

So a proof by contradiction would start out as: Assume \(r\) is rational, and \(r+s\) is not rational, and \(s\) is rational. We play with this until we reach something that is false.

OpenStudy (anonymous):

Okay, so we only negated the conclusion, gotcha. So do we still have to work in the direction of if -> then, or do we just want any sort of contradiction?

OpenStudy (anonymous):

At this point, we want any contradiction. Using those three statements.

OpenStudy (anonymous):

So if r is rational, it can be written as something like m/n and if s is rational, it can be written as something like p/q. So r + s is:\[\frac{ m }{ n }+ \frac{ p }{ q }\]But...what kind of expression do we have to represent irrational? Like, I can say rational is an expression m/n, but what is irrational?

OpenStudy (freckles):

m,q,n, and p are what kinda numbers?

OpenStudy (freckles):

And what happens when you combine those fractions?

OpenStudy (anonymous):

Oh, integers where n and q are nonzero.

OpenStudy (anonymous):

\[\frac{ m }{ n }+ \frac{ p}{ q }= \frac{ mq + np }{ nq }\]

OpenStudy (freckles):

is mq+np an integer? is nq an integer?

OpenStudy (anonymous):

Based on closure, I would think so...?

OpenStudy (freckles):

yeah but that would mean r+s is what kinda number?

OpenStudy (anonymous):

Right, so r+s is an integer, but that's a contradiction. Alright, cool. So the only problem is this analysis class is REALLY picky about using only what we have been given. So I'm not sure whether or not I could make a closure argument. Is there another way it might be able to be stated?

OpenStudy (freckles):

r+s is a rational, but that's a contradiction*

OpenStudy (anonymous):

Oops, lol.

ganeshie8 (ganeshie8):

looks we gave up cheaply on proving the contrapositive hmm

OpenStudy (anonymous):

I wouldn't know which way is best, lol. But yeah, is there another way to come to the above result, that r+s is rational without making a closure argument? Because we haven't explicitly discussed closure.

ganeshie8 (ganeshie8):

i think contradiction and contrapositive take same effort

OpenStudy (freckles):

I didn't know how to do the contrapostive way

OpenStudy (anonymous):

The contrapostive isn't a bad idea, it just sounds a little strange in my opinion.

ganeshie8 (ganeshie8):

oh you don't like saying \(\large (n,q) \in \mathbb Q \implies nq \in \mathbb Q\)

OpenStudy (freckles):

If s is rational, then r is not rational or (r+s) is rational. We get to suppose s is rational...But r isn't in the hypothesis so I don't know what to do with that. So I don't see a direct constrapositve proof

OpenStudy (freckles):

could be wrong

OpenStudy (anonymous):

There is no way to do this proof without knowing the fact that the sum of two rational numbers is still a rational number. That is what we mean by closure. You don't have to specifically say "closure", but you do need that the sum of two rational numbers is a rational number.

ganeshie8 (ganeshie8):

^^

OpenStudy (anonymous):

The issue that freckles has with the contrapositive is basically my qualm as well. It just sounds...strange haha.

OpenStudy (anonymous):

It does work though. It would go like this. Assume s is rational. We want to show that either r is irrational, or r+s is rational. If r is irrational, we are done. If not, then r is rational, and since s is rational, r + s is rational..

OpenStudy (anonymous):

See how it just kind of sounds....weird? Its totally equivalent to a direct proof or a a proof by contradiction, but it just has a strange vibe to it O.o

ganeshie8 (ganeshie8):

nice :) the logical disjunction arguemnt... looks more of a proof in logic class !

OpenStudy (freckles):

That is neat.

OpenStudy (anonymous):

Ah, okay, so that's how it would work with the contrapositive. Thanks very much, all of you, I really appreciate it ^_^

ganeshie8 (ganeshie8):

so if i were dirac, i would call contradiction proof an elegant one and the contrapositive proof a weirdo ;p

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

I would agree :) Although one of my professors would yell at me for doing too many proofs by contradiction >.< He said that real men do things directly >.>

OpenStudy (freckles):

weirdos need to exist where would we be without Einstein? i heard he was weird

ganeshie8 (ganeshie8):

lol i thought contrapositive is also a kind of indirect proof as we are switching around the statements and negating.. i think its more indirect than a simple contradiction.. i dont know what your prof would call a man doing all contrapositive proofs hmm

OpenStudy (freckles):

what if you want to do the a proof by contradiction of the contraspotive would that be a indirectest proof

OpenStudy (freckles):

i'm being dumb you don't have to answer that

OpenStudy (anonymous):

haha, thats awesome.

ganeshie8 (ganeshie8):

thats a frecky idea haha! it should work atleast logically a proof by contradiction of a contraapositive of a given statement will be a real turn on for a logic prof ;)

OpenStudy (freckles):

We would have if p then q. Then turn it into not q then not p then by contradiction we would assume p and not q

OpenStudy (freckles):

but if wanted to do if p then q by contradiction then we would assume not q and p it would be logically equivalent for sure

ganeshie8 (ganeshie8):

Woah! that looks like a round trip, thats all i guess... its same as the contradiction of original statement's conclusion

OpenStudy (freckles):

right

OpenStudy (freckles):

it is a circle

OpenStudy (freckles):

that would be a question on a test.. i don't know how to phrase it exactly

OpenStudy (freckles):

good question*

OpenStudy (freckles):

Is it pointless to prove some thing by contradiction of the contrapositive?

OpenStudy (freckles):

That is the only way I can think of how to phrase it

ganeshie8 (ganeshie8):

lol im still thinking... for p->q both the statemetns in `direct contradiction` and `contradiction of contrapositive` are same : \(\large p \land \neg q\)

ganeshie8 (ganeshie8):

so `contradiction of contrapositive` may not ease out things that the original `direct contradiction` cannot... as the statement we start with in both proofs are same hmm

OpenStudy (anonymous):

@freckles I apologize to call you back. I thought the discussion might give me an idea on the 2nd part of the problem, but it wasn't something that jumped to me. I figured it would be best to take the same approach, do negation and say s is rational. In the first part, we were able to combine fractions and make a rational expression. But it seems like the second problem would want us to construct an irrational expression. But I'm not sure what such an expression would look like. Any ideas?

OpenStudy (freckles):

try contradiction again

OpenStudy (anonymous):

Yes, that is where I was going. The contradiction would be to assume s is rational though, correct? Since that is the conclusion. But I was looking at it and trying to see which of the statements I'd be able to somehow contradict. I'd have to somehow get either r to be rational, s = 0, or rs to be irrational. I don't know what an irrational expression would even look like. Yet I felt like somehow I would have to construct such a statement to get a contradiction.

OpenStudy (freckles):

\[s \text{ is rational } , rs \text{ is rational}, s \neq 0, \text{ and } r \text{ is not rational} \]

OpenStudy (freckles):

\[\text{ Let } a,b,c,d \in \mathbb{Z} \text{ with } b,d \neq 0 . \\ \text{ Let } s=\frac{a}{b} \text{ and } rs=\frac{c}{d}\]

OpenStudy (freckles):

\[rs=r \frac{a}{b}=\frac{c}{d}\] Ans since s does not equal 0 then a does not equal 0 so \[r=\frac{c}{d} \frac{1}{s}=\frac{c}{d} \frac{b}{a}\]

OpenStudy (freckles):

what do you notice here?

OpenStudy (anonymous):

Im not sure, lol. Like, I'm not sure if there is something funky about either of those expressions. I see the equality and how you got them, just not sure what is off about them immediately.

OpenStudy (anonymous):

Oh, does that make R automatically rational?

OpenStudy (anonymous):

r*

OpenStudy (freckles):

yes we have a product of integers on top and on bottom which would make r rational

OpenStudy (anonymous):

Lol, thats actually kinda funny, I spent the whole time looking at things similar to that wondering how I was going to get something that looked irrational (to contradict rs is rational) when all I had to do is see that r was rational with those manipulations x_x

OpenStudy (freckles):

Sometimes I look at things forever too and wonder why they aren't making sense and then i give up and sometimes post and ganiesh always knows what to say.

OpenStudy (freckles):

or whatever his name is

OpenStudy (anonymous):

Yeah. I mean, there's always logic to these that can be seen, I'm just not used to working with them and manipulating them yet. And if I also may ask, when you and ganeshie were talking earlier, ganeshie mentioned logical disjunction argument...what is that by chance?

OpenStudy (freckles):

That is kinda a term I forgot and would have to look up

OpenStudy (freckles):

But he was referring to nerd's post

OpenStudy (freckles):

if i remember correctly

OpenStudy (anonymous):

Yeah. I just hadn't heard of it. I've heard of disjoint, but not sure if thats the same.

OpenStudy (freckles):

Assume s is rational. We want to show that either r is irrational, or r+s is rational. If r is irrational, we are done. If not, then r is rational, and since s is rational, r + s is rational..

ganeshie8 (ganeshie8):

disjunction is a fancy name for "logical or"

ganeshie8 (ganeshie8):

lets mimic nerd's proof by contrapositive for part 2 also maybe ?

OpenStudy (anonymous):

I wouldnt mind seeing how that works ^_^

ganeshie8 (ganeshie8):

statement : r is not rational, s is not equal to 0, and rs is rational implies s is not rational contrapositive : s is rational implies "r is rational" or "s = 0" or "rs is not rational"

OpenStudy (anonymous):

Yeah, had no idea how you'd state the contrapositive with 3 statements like that. Had no idea if you'd make them all "or" statements or not :P

ganeshie8 (ganeshie8):

if r is rational or s = 0, we are done. if not, rs = irrational*s = irrational \(\large \square \)

ganeshie8 (ganeshie8):

yes pack those 3 statements before implies as one : ` r is not rational, s is not equal to 0, and rs is rational` implies s is not rational

ganeshie8 (ganeshie8):

negation of AND of several several statements is OR of negation of each statemetn : \[\large \neg(A\land B\land \cdots ) = \neg A \lor \neg B \lor \cdots \]

OpenStudy (freckles):

not(A and B) =notA or notB so if B=(C and D) =notA or not(C and D) =notA or (notC or notD) =notA or notC or notD

ganeshie8 (ganeshie8):

^^

OpenStudy (anonymous):

That contrapositive seems....off. Like, it doesnt seem like youre doing any sort of proof. I guess the 2nd part of the statement is using the idea that irrational * rational = irrational (I know it's true, but is that a some sort of actuak explicit property or is it a given?). But for the first part, yeah, if r is rational or s is 0 then it's true, but you don't have to really show anything for that?

ganeshie8 (ganeshie8):

we are translating the given problem into an equivalent statement

ganeshie8 (ganeshie8):

contrapositive is logically equivalent to given statement, so proving a contrapositive is same as proving the given statemetn. so yes, the actuall proof starts only after taking the contrapositive